# Infrared Object Detection with Ultra Small ConvNets: Is ImageNet Pretraining Still Useful?

### v1

@@ Abstract
Many real-world applications require recognition models that are robust to different operational conditions and modalities, but at the same time run on small embedded devices, with limited hardware. While for normal size models, pre-training is known to be very beneficial in accuracy and robustness, for small models, that can be employed for embedded and edge devices, its effect is not clear. In this work, we investigate the effect of ImageNet pretraining on increasingly small backbone architectures (ultra-small models, with 1M parameters) with respect to robustness in downstream object detection tasks in the infrared visual modality. Using scaling laws derived from standard object recognition architectures, we construct two ultra-small backbone families and systematically study their performance. Our experiments111The code will be available online after acceptance. on three different datasets reveal that while ImageNet pre-training is still useful, beyond a certain capacity threshold, it offers diminishing returns in terms of out-of-distribution detection robustness. Therefore, we advise practitioners to still use pre-training and, when possible avoid too small models as while they might work well for in-domain problems, they are brittle when working conditions are different.
@@ 1 Introduction
Visual recognition is a fundamental tool for many different tasks such as autonomous driving, robotics, security, medical and satellite imaging [19, 15, 22, 4]. Two features are essential for the effective use of visual recognition in real-world applications: i) the ability to generalize to different environments, modalities and conditions ii) the capability to run on embedded devices.
While there is a large body of research in robustness [29] and small recognition models [39, 17], the simultaneous use of the two has been mostly overlooked. In this research, we focus on robust models that can adapt to different situations and contexts but can also run on low budget embedded devices. The conjunction of the two features makes the task harder, as a typical way to increase the robustness is to use a larger model, with more parameters and computational cost [27], which is incompatible with the reduced resources of low budget embedded devices.
Commercial GPU Mobile Tiny Ultra-Small (NVIDIA RTX) (iPhone 16) (Cortex-A53) (Cortex-M7) Models B7 to B4 B4 to B1 B0 to B-2 B-3 to B-7 FLOPs 82T 2.15T 12.3G 2.0G Memory 24GB 8GB 2GB 1GB
Since the advent of deep learning, one key approach to add robustness to models is to use a pre-training phase [9, 8], in which a model is trained on a large body of data, different than the final task. With large models, this approach helps to bootstrap the model with useful and general features, which will then be refined in a second phase of fine-tuning, on a smaller dataset on the specific task and data distribution. This approach is powerful because the expensive pre-training can be done once, and then the same model is fine-tuned for many different tasks. In addition, even when the data distribution and the task of the final model are different from the pre-training data, the final model still gains form it [18, 44]. This gain is mostly because deep learning models are trained with a non convex objective and they have millions of parameters to learn. Thus, during training, the optimization can find multiple local minima solutions. With pre-training, we make sure the fine-tuining phase starts from a good initialization in which the model has already seen many data points, such that the final solution will be better and more robust [11].
While pre-training is well-known to help for large deep-learning models, we also know that for linear models and common losses (e.g. cross-entropy), pre-training is not needed [1] because the optimization is convex and we will obtain always the same solution, independently from the initialization. In this work we analyse the effect of pre-training when using very small, but still non-linear models (ultra-small), in which the number of parameters is very low (M). Table 1 provide a summary of different ranges of computational resources and memory needed for different kind of applications and its associated models in the extended EfficientNet family. As normally the number of samples should be somewhere proportional to the number of data-points [16], we expect that when the number of parameters is small enough, pre-training will stop to be useful. Thus, in this work, we characterize the effect of pre-training for performance generalization when reducing the model size. We believe that this is an important questions that has been overlooked by the vision community. In fact, by knowing in which situation and how much pre-training helps a small model to better generalize, it can help practitioners and industrial design to make a more appropriate choice in the trade-off between model size, performance and generalization.
More in detail, this study focus on object detection, which is one of the most fundamental tasks in computer vision [9, 32, 8], because it allows the model to not only recognize and object, but also find its position in an image, which is very important for certain tasks, such as autonomous driving or robotics. Furthermore, we specialize our research to infrared data, as it is one of the most important modalities on embedded device, as it allows night vision while keeping most of the visual information needed for image understanding [36].
The main contributions of our work can be summarized as follows:
(1) We analyze for the first time the effect of RGB images (ImageNet) pre-training with ultra small infrared object detection models.
(2) We evaluate two well-known families of convolutional models (EfficientNet and MobileNet), on three different infrared datasets, by scaling down the size of the models and testing their resilience to generalization with or without pre-training.
(3) With the support of important ablations on the amount of data and resolution for the fine-tuning phase, our empirical evidence show that pre-training is important for the generalization of very small models, but when the size becomes too small (100k params), pre-training has negative or no effect on generalization performance.
@@ 2 Related Work
Object Detection. Unlike classification, which focuses only on identifying object categories, object detection requires both classification and precise localization within an image. Deep learning-based detectors are typically categorized into two-stage and one-stage methods. Two-stage detectors, such as Faster R-CNN [34], employ a Region Proposal Network (RPN) to generate candidate regions, enabling accurate detection with reasonable efficiency. One-stage detectors, like YOLO [33] and FCOS [41], eliminate the proposal step, resulting in faster inference. More recently, detectors incorporate language in their components, such as Grounding DINO [24] and YOLO-World [5]. Although this leads to an increase in performance, the efficiency is compromised, especially for edge devices. Therefore, different detectors suitable for edge devices have been explored, for instance, NanoDet [31], which proposes a FCOS-style one-stage anchor-free detection model capable of having a good tradeoff between inference speed and memory for hardware-constrained edge devices.
Pretraining for Object detection. Pretraining is a common practice in object detection, significantly boosting performance, especially for larger models [9, 32]. It has been shown to accelerate convergence and enhance accuracy when transferring to detection tasks with limited training data [9, 18]. Traditionally, ImageNet pretraining has been the standard, providing robust features that transfer well to various downstream tasks. Recent studies have explored self-supervised pretraining [3, 12] to further improve downstream performance. However, these works predominantly focus on large or standard-scale models. The impact of pretraining on ultra-small models, particularly those with backbone size up to two orders of magnitude smaller than EfficientNet-B0 [39], remains underexplored. Additionally, the effect of pretraining on generalizing to out-of-domain detection tasks is less studied. In this work, we analyze supervised ImageNet pretraining for ultra-small convolutional networks in out-of-domain infrared object detection settings, highlighting the benefits and limitations in resource-constrained scenarios.
Out-of-distribution Robustness.
Robustness to distribution shifts remains a critical challenge in object detection, particularly in safety-critical applications. Prior works have studied domain adaptation [37], data augmentation [7, 45], and self-training [48] to improve generalization across domains. Recent efforts focus on detecting out-of-distribution images [43] and improving model calibration under distribution shift [30]. While these approaches typically employ large backbone networks to enhance robustness, the behavior of ultrasmall models under OOD settings is underexplored. Our work investigates the OOD robustness of object detection models that are orders of magnitude smaller than conventional architectures, addressing the need for efficient yet robust detectors in resource-constrained deployment scenarios.
@@ 3.1 Preliminaries
Object Detection.
The training set for OD is denoted as , where represents an image in the dataset, with dimensions and channels. Subsequently, the OD model aims to identify regions of interest within these images, denoted as . The top-left corner coordinates and the width and height of the object define each region of interest . Additionally, a classification label is assigned to each detected object, indicating its corresponding class within the dataset. In terms of optimization, the primary goal of this task is to maximize detection accuracy, often measured using the average precision (AP) metric across all classes. An OD is formally represented as the mapping , where denotes the parameter vector. To effectively train a detector, a differentiable surrogate for the AP metric, referred to as the detection cost function, , is employed. The typical structure of such a cost function involves computing the average detection loss over dataset , denoted as , described as:
@@
We systematically construct ultra-small object detectors for non-RGB modalities by downscaling two backbone networks, EfficientNet-B0 [39] and MobileNetV3-small [13]. Specifically, we reduce both the width and depth of these models up to seven levels in case of EfficientNet down EfficientNet-B0 and six levels down MobileNet-V3 small while maintaining input resolution, which is critical for object detection performance. Details of our downscaling procedure are provided in the supplementary material.
3.2 Proposed robustness evaluation pipeline
@@
Figure 2 presents an overview of our framework. The pipeline comprises three main steps: (1) initializing the detector backbone either with ImageNet pre-training or random initialization, (2) training the detector on in-domain data, and (3) evaluating out-of-domain robustness to compare the effectiveness of ImageNet pre-training against random initialization.
Supervised ImageNet Pre-training.
Supervised pre-training on ImageNet [6] has been a standard approach to initialize convolutional neural networks for various downstream tasks. In this setting, models are trained on the ImageNet-1K dataset, which contains approximately 1.28 million training images across 1000 object categories. The objective is to minimize the cross-entropy loss between the predicted class probabilities and the ground truth labels:
where is the number of classes, is the one-hot encoded ground truth label, and is the predicted softmax probability for class .
The pre-training process results in models with weights that capture generic visual representations, which are transferable to downstream tasks such as object detection, including adaptation to different modalities. We adopt this standard supervised pre-training to initialize our Object Detector backbones to compare them against randomly initialized backbone weights.
Fine-tuning Object Detectors.
@@
We study supervised object detection using data from an in-domain distribution , where denotes an image with spatial dimensions and channels. Our goal is to evaluate the robustness of object detectors when tested on an out-of-domain distribution , comparing models initialized with random weights versus Imagenet pre-trained weights, and examining how model size affects this robustness. We train two versions of each detector:
@@
Random initialization: All layers are initialized from scratch without any pre-training.
ImageNet pre-training: Backbone networks are initialized with Imagenet pre-trained weights as described in Section 3.2, while the rest of the detector is randomly initialized.
@@
Both variants are trained using supervised in-domain data following the procedure in Section 3.1. This ensures that the detectors learn representations useful for both in-domain and OOD Object Detection.
Robustness Evaluation.
@@
We assess the relationship between model size and robustness of Imagenet pre-training on two different ways. First, we assess modality robustness where a detector is fine-tuned on RGB images and then evaluated on infrared data. The second evaluation involves a dataset with multiple cameras/views. In this case, the fine-tuning is performed on some viewpoints, while the evaluations is run on other viewpoints, not seen during fine-tuning.
@@ 4 Experiments
| Model | Backbone Params | Backbone MFLOPs | Detector Params | Detector MFLOPs |

### v2

# Pretraining Helps When Capacity Allows: Evidence from Ultra‑Small ConvNets

@@ Abstract
Robust visual recognition on embedded platforms requires models that both generalize out‑of‑distribution (OOD) and fit into tiny compute/memory budgets. While pre‑training is a standard route to robustness for mid/large backbones, its value in the ultra‑small regime remains unclear. We present a capacity‑aware study of pre‑training for two efficient ConvNet families (EfficientNet and MobileNetV3) scaled from “small” to “ultra‑small” via a simple, reproducible recipe. We compare three initializations — ImageNetCOCO pretraining, ImageNet classification pretraining, and training from scratch—on two axes of distribution shift: (i) cross‑dataset RGBRGB transfer between LLVIP and FLIR (ii) cross‑modality detection where models are fine‑tuned on RGB and evaluated on infrared (IR). A complementary classification study on DomainNet probes whether the trends extend beyond detection. Across settings, we find that pretraining’s benefit is conditional on both backbone capacity and shift difficulty. Task‑aligned ImagenetCOCO pretraining is the most reliable starting point at moderate sizes and for the easier transfer direction. In the low‑capacity regimes, differences are typically within run‑to‑run variation, and training from scratch can match or surpass pre‑training. Classification mirrors this capacity gating. Our results test the premise ”pretraining always helps” and instead quantify when task‑aligned pretraining pays off for ultra‑small backbones and when it likely does not111https://github.com/srikanth-sfu/wacv26-ultrasmall-cnns.
@@ 1 Introduction
Visual recognition supports applications from autonomous driving and robotics to security, medical imaging, and remote sensing [19, 15, 23, 5]. Two properties are crucial for deployment: generalization across environments, sensors, and conditions, and efficiency, enabling real-time inference on embedded devices. While generalization [28] and efficiency [37, 17] have been widely studied, their joint pursuit—generalizable yet ultra-small models remains underexplored. Generalization is often achieved through large-scale pretraining followed by fine-tuning [9, 8], which reliably benefits high-capacity models [18, 41, 11]. At the other extreme, classical convex settings suggest little dependence on initialization [1]. The poorly understood regime lies in between: how much does pretraining help when shrinking modern vision backbones to the ultra-small (sub-million-parameter) scale?
We examine this question with two efficient ConvNet families (EfficientNet, MobileNetV3), three initializations (COCO, ImageNet, scratch), and two shift types: cross-modality (RGBIR) and cross-dataset (RGBRGB). Our main application is object detection, motivated by embedded night-time perception [34], complemented by domain generalization for classification to assess generality of our findings beyond our principal application. To densify the capacity axis, we extend EfficientNet (B7B0) with BB and MobileNetV3 with SS (Table 1).
Our study shows that pretraining benefits are conditional, not universal. For cross-dataset generalization for object detection ImageNet-COCO pretraining (also referred to as INCOCO pretraining) usually outperforms ImageNet pretraining (also referred to as IN) and scratch, but effects shrink with capacity and can vanish in harder directions (e.g., LLVIPFLIR). For cross-domain generalization INCOCO is most reliable, especially on FLIR, though scratch competes at the smallest sizes. For classification, IN improves large to moderate sized models, but gains fade or reverse for smallest of model sizes. Our additional controlled experiments on impact of fine-tuning data size confirm the same capacity-dependent patterns (Fig. 1). Our findings differ from common notion that supervised pretraining recipes are necessary to obtain better generalization regardless of the model properties. Benefits depend jointly on model capacity and shift difficulty: clear gains appear for moderate sizes or easier shifts, but drop to noise below a certain threshold for difficult transfers.
Our contributions can be summarized as follows:
A capacity-aware evaluation of EfficientNet and MobileNetV3 across detection and classification, with INCOCO, IN, and scratch initializations.
New ultra-small variants (BB, SS) and an explicit scaling recipe for reproducibility.
Empirical evidence that pretraining benefits diminish with extreme capacity reduction, sometimes making random initialization competitive.
Practical guidance for embedded deployment: when pretraining is worthwhile versus when task-specific data tuning matter more.
Our aim is not to propose a new backbone but to quantify capacity - pretraining interactions under embedded constraints, providing actionable evidence for design at the low-capacity frontier.
| Commercial GPU | Mobile | Tiny | Ultra-Small | |
| (NVIDIA RTX) | (iPhone 16) | (Cortex-A53) | (Cortex-M7) | |
| Models | B7 to B4 | B4 to B1 | B0 to B-2 | B-3 to B-7 |
| FLOPs | 82T | 2.15T | 12.3G | 2.0G |
| Memory | 24GB | 8GB | 2GB | 1GB |
@@ 2 Related Work
Out-of-distribution Robustness.
Robustness to distribution shifts remains a critical challenge in object detection, particularly in safety-critical applications. Prior works have studied domain adaptation [35], data augmentation [7, 42], and self-training [45] to improve generalization across domains. Recent efforts focus on detecting out-of-distribution images [40] and improving model calibration under distribution shift [29].
Pretraining is a common practice in tasks such as image classification and object detection, significantly boosting performance, especially for larger models [9, 32]. It has been shown to accelerate convergence and enhance accuracy when transferring to tasks with limited training data [9, 18]. Traditionally, ImageNet pretraining has been the standard, providing robust features that transfer well to various downstream tasks. Recent studies have explored self-supervised pretraining [3, 12] to further improve downstream performance. However, these works predominantly focus on large or standard-scale models. The impact of pretraining on ultra-small models, particularly those with backbone size up to two orders of magnitude smaller than EfficientNet-B0 [37], remains underexplored. Our work investigates the OOD robustness of models that are orders of magnitude smaller than conventional architectures, addressing the need for efficient yet robust frameworks in resource-constrained deployment scenarios.
@@ 3.1 Preliminaries
We systematically construct ultra-small models by downscaling two backbone networks, EfficientNet-B0 [37] and MobileNetV3-small [13]. Specifically, we reduce both the width and depth of these models up to seven levels in case of EfficientNet down EfficientNet-B0 and six levels down MobileNet-V3 small while maintaining input resolution, which is critical for downstream tasks such as object detection. Details of our downscaling procedure are provided in the supplementary material.
3.2 Proposed generalization evaluation pipeline
@@
Figure 2 presents an overview of our framework. The pipeline comprises three main steps: (1) initializing the detector backbone either with ImageNet pretraining or with an additional optional task specific supervised finetuning such as with COCO dataset for object detection or random initialization, (2) finetuning models on in-domain data, and (3) evaluating out-of-domain generalization to compare the effectiveness of different pretraining against random initialization.
Supervised ImageNet pretraining (IN)
Supervised pretraining on ImageNet [6] has been a standard approach to initialize convolutional neural networks for various downstream tasks. In this setting, models are trained on the ImageNet-1K dataset, which contains approximately 1.28 million training images across 1000 object categories with standard recipe of using cross entropy loss. The pretraining process results in models with weights that capture generic visual representations, which are transferable to downstream tasks and domains. We adopt this standard supervised pretraining to initialize our backbones and compare them against no pretraining as well as task-specific pretraining (when applicable) for mutliple tasks including object detection and Image classification for various scales of ultra-small models. We refer to this pretraining paradigm as IN in subsequent sections, figures, and captions.
Supervised COCO finetuning(INCOCO)
For object detection task, we further also consider supervised fine-tuning on the COCO detection dataset, after standard ImageNet-1K supervised backbone initialization [22]. In this setting, models are trained on the COCO dataset with the NanoDet standard recipe [31]. The pretraining process results in models with weights that capture detection-specific representations, which are transferable to downstream object detection tasks that generalize across different domains and modalities. In this work, we assess this pretraining benefit for different levels of small models against task-agnostic pretraining and no pretraining variants. We refer to this two-stage detection-aligned pretraining paradigm as INCOCO in all subsequent sections, figures, and captions.
Finetuning Object Detectors.
@@
We study supervised object detection using data from an in-domain distribution , where denotes an image with spatial dimensions and channels. Our goal is to evaluate the generalization capability of object detectors when tested on multiple generalization tasks , comparing models initialized with random weights versus Imagenet pretrained weights, and examining how model size affects different types of generalization. We train three versions of each detector:
@@
Random initialization: All layers are initialized from scratch without any pretraining.
Task Independent large scale pretraining: Backbone networks are initialized with IN pretrained weights.
Task-specific pretraining: Backbones are initialized with task-specific pretraining paradigm, such as INCOCO for object detection, and are used to initialize most of the detector layers.
@@
All the variants are trained using supervised in-domain data. This ensures that the detectors learn representations useful for both in-domain and generalization settings.
Generalization Evaluation.
@@
We assess the relationship between model size and pretraining generalization from multiple perspectives. First, we assess generalization cross-domain(s) for both object detection and image classification. Second, we test cross-modal generalization where a detector is fine-tuned on RGB images and then evaluated on infrared data. The third evaluation protocol involves viewpoint generalization (using our In-house Distech dataset) with multiple cameras/views. In this case, the fine-tuning is performed on some viewpoints, while the evaluations are run on viewpoints not seen during fine-tuning.
@@ 4 Experiments
| Model | Backbone | Detector | ||
| Params | MFLOPs | Params | MFLOPs | |
[…]